OpenStudy (brittany):

solve using the quadtratic formula 2x^2=4-7x

OpenStudy (anonymous):

\[2x^2+7x-4=0\] now use quadratic formula

OpenStudy (brittany):

im not sure how to

OpenStudy (anonymous):

how to ???

OpenStudy (brittany):

how to make it into quadratic formua

OpenStudy (anonymous):

\[ax^2+bx+c=0\] \[x = (-b \pm \sqrt{b^2-4ac})/2a\] where a=2,b=7 and c=-4

OpenStudy (anonymous):

just plug in the values